The pressure P,volume V,absolute temperature T of a real gas are related by the equation (P+a/V2)(V-b)=RT.Here a,b,r are constants .The dimension of the constant a is

Grade:

4 Answers

Ritvik Gautam
13 years ago

We know that only those quantities can be added or subtracted which have the same dimensions.

Here, P and a/V2 are added which means they have the same dimensions.

Therefore,

[M1 L-1 T-2 ] = a/[M0 L 6 T 0 ]

a = [M1 L-1 T-2 ] x [M0 L 6 T 0 ]

a = [M1 L5 T-2 ]

 

 

Please Approve!!

aryan
5 years ago
acc to law of dimensional homogenity , the dimensions on the left hand side should be equal to the dimensions on the right hand side . therefore, RT = P = A/V^2 . We know that only those quantities can be added or subtracted which have the same dimensions.

Here, P and a/V2 are added which means they have the same dimensions.

p= a/v^2 
ml^-1 t^-2 = a / m^0 l^6 t^0 
a = M^1 L^5 T^-2
Navya vats
5 years ago
Can have simple trick, 
Hello ......Here navya ,
Your solution 👇
We know that only same quant.  Can be added or substracted so 👇✌
We have (p+a/ve) = a=pv2 = ML-1T-2 × (L3 )2
= ML-1T-2 × L6 ={ ML5T2}👍😊
Navya vats
5 years ago
Upper there is ML5T-2 ✌✋👍 so fr mistake .....Make sure to correct this only is correct bye .    
.....
. ......
.........
 
 
 
 
 
 
 
 
.............
..
 
 
 
........